OpenStudy (anonymous):

I'm having a little trouble figuring this out, can someone help? Melissa bought a new car with 12 optional features to the car. Each optional feature costs either $23.99 or $120.84. If she paid $675.28 for all the optional features, how many of each type did she buy? Write and solve an equation to solve this problem. Show all your work.

OpenStudy (solomonzelman):

Okay, let: Number of the purchases first feature \(\normalsize\color{royalblue}{ \rm F_1 }\) Number of the purchases second feature \(\normalsize\color{royalblue}{ \rm F_2 }\) \(\normalsize\color{royalblue}{ \rm F_1+F_2 =12 }\) \(\normalsize\color{royalblue}{ \rm (23.99)F_1+(120.84)F_2 =120.84 }\)

OpenStudy (solomonzelman):

mstake

OpenStudy (solomonzelman):

Your system should look like this. \(\normalsize\color{royalblue}{ \rm F_1+F_2 =120.84 }\) \(\normalsize\color{royalblue}{ \rm (23.99)F_1+(120.84)F_2 =675.28 }\)

OpenStudy (solomonzelman):

you can use \(\normalsize\color{royalblue}{ \rm x }\) for \(\normalsize\color{royalblue}{ \rm F_1 }\) AND \(\normalsize\color{royalblue}{ \rm y }\) for \(\normalsize\color{royalblue}{ \rm F_2 }\)

OpenStudy (anonymous):

@SolomonZelman Thank-You!

OpenStudy (solomonzelman):

yw

OpenStudy (mathmate):

@SolomonZelman I think you mean: \(\normalsize\color{royalblue}{ \rm F_1+F_2 =12 }\)

OpenStudy (solomonzelman):

yes, that too... it's a day of mistakes (as always)

OpenStudy (solomonzelman):

this is the system: \(\normalsize\color{royalblue}{ \rm F_1+F_2=12 \color{white}{\Huge|} }\) \(\normalsize\color{royalblue}{ \rm(23.99)F_1+(120.84)F_2=120.84 }\)

sammixboo (sammixboo):

Hello and welcome to OpenStudy @Zoe9639 :) Since your question has been answered, you should close it so other users know that it has been answered! :) Do you know how to close your question? Also, you can't post another question in this subject until you have closed this one.

OpenStudy (anonymous):

@sammixboo thank you! I forgot :)

sammixboo (sammixboo):

No problem, and it's okay!